# PLeaS — Merging Models with Permutations and Least Squares

### v1

@@ 4.2 Permuted least squares
Suppose, for example, that the target merged model has the same architecture and size as each of the base models. Once the permutations, $P_{i}$’s, have been computed, we propose optimizing the weight matrices of the merged model by solving the following least-squares problem:
$W_{i}^{M}\;\in\;\operatorname*{arg\min}_{W\in\mathbb{R}^{d_{i}\times d_{i+1}}}\|(Z_{i}^{A}+P_{i}Z_{i}^{B})W-(Z_{i+1}^{A}+P_{i+1}Z_{i+1}^{B})\|^{2}\;,$ (1)
independently for each layer $i\in[L]$. This is motivated by the impressive performance of the ensemble method (e.g., [33] and Section 5), which retains two separate models and only averages the (permuted) activations at the last layer (before softmax): $\tilde{z}_{L+1}=z^{A}_{L+1}+z^{B}_{L+1}$. We aim to have our merged model approximate such activations. We inductively assume that the first $i-1$ layers are properly merged. Hence, the ensemble of the permuted features (of the $i^{\rm th}$ layer) of the component models can be well approximated by the activations at the input of the $i^{\rm th}$ layer of the merged model. We denote the ensembled features by $\tilde{Z}_{i}=(Z_{i}^{A}+P_{i}Z_{i}^{B})\in{\mathbb{R}}^{d_{i}\times n}$. The goal of the above optimization is to match the ensembled activation of the next layer, $\tilde{Z}_{i+1}=(Z_{i+1}^{A}+P_{i+1}Z_{i+1}^{B})$, with a linear transform of the input ensemble: $\tilde{Z}_{i}W$. We empirically validate this choice of using a permuted ensemble of features to optimize the weights of the merged model in Table 3 in Section B.1, where we compare with alternatives choices to Eq. (1).
This second step of PLeaS is similar to feature distillation. However, the key novelty arises from averaging the permuted features for transferring knowledge from multiple models. This is critical for accurate prediction. In Section 5, we compare PLeaS against RegMean [14], which optimizes an objective similar to Eq. 1 without the permutations and averaging. This method merges models by minimizing $\|Z_{i}^{A}W-Z_{i+1}^{A}\|^{2}+\|Z_{i}^{B}W-Z_{i+1}^{B}\|^{2}$. As we show in Section 5, RegMean performs poorly compared to PLeaS . Apart from the inference computation budget for the final model, PLeaS is completely hyperparameter free.
Note that the second step of PLeaS is fully compatible with the partial merging of Section 4.1 as well: we can directly set the values of $W_{i}^{M}$ corresponding to the unmerged features to be the respective values of $W_{i}^{A}$ and $W_{i}^{B}$. While the objective in Eq. 1 can be minimized in closed form using Ordinary Least Squares (OLS), we practically implement it using gradient descent for ease of use with convolution layers. Given that the objective is convex if computed layer-wise, the weight matrices $W_{i}^{M}$ converge in relatively few steps (less than 100) of gradient descent. Further, we solve this optimization independently for each layer, which can be efficiently parallelized.
@@ 4.3 Data requirements of PLeaS
PLeaS has two steps. The first step finds permutations to match features using weight or activation matching. The second step computes weight matrices to mimic the ensemble of the merged features more closely. In order to compute these features, one could use the data from the training domains, however, this may not be feasible for privacy or commercial reasons. Hence, we propose an alternative scheme—dubbed PLeaS-free —which uses a general vision dataset, like ImageNet, to compute the activations of the component models. These activations are then used to merge domain specific models without requiring any data from their training domains. In Section 5.4, we demonstrate that PLeaS-free suffers minimal degradation compared to PLeaS, suggesting a wider applicability.
@@ 5.1 Settings
We show the effectiveness of our method in merging ResNet-50 models fine-tuned on different datasets starting with different initializations. We consider the following experimental settings:
  • Section 5.2 and Section 5.3 deal with the problem of merging models fine-tuned on different datasets with shared or different label spaces respectively.
  • Section 5.4 considers the setting where we do not have access to data from the training domains.
  • Section 5.5 deals with merging models which were fine-tuned from the same base model.
  • Finally, Section B.4 considers a setting where we want to merge differently initialized models trained on the same dataset, and Section B.1 provides an ablation study over the choice of loss functions.
Unless specified otherwise, we use the weight matching strategy from [1] to compute the permutations for PLeaS . For each task, we also report results for Permutations, which is the model obtained by weight averaging the component models after applying the permutations obtained from the first step of PLeaS . Following the recommendation from REPAIR [15], we recompute the batch-norm parameters of the model after merging for all methods. We run each merging experiment for three different seeds, and across two different initial models. We find that inter-run deviation in performance is low, with the standard deviation usually being less than 1%. We report disaggregated results along with these standard deviations in Section B.5.
Baselines
We compare our method against prior works including Git Re-Basin[1], Simple Averaging [12], RegMean [14] and ZipIt! [28]. We also consider two practical upper bounds – training a router model based Mixture of Experts model (MoE), and ensembling the predictions (or activations) of the original models. The former requires storing both models and running one of them based on the router, hence having $\times1$ FLOPs and 2×2$\times$2 × memory requirements, while the latter requires running both the models in parallel, and hence has 2×2$\times$2 × FLOPs and memory requirements. We find that the performance of the ensemble and MoE models is close to the best performance of a single model on the dataset that it was trained on.
5.2 Models fine-tuned with a shared label space
Table 1: Merging models trained on pairs of different datasets: For each pair of datasets, we merge models and compute the final accuracy on the pair. To compute the final accuracy on a dataset, we average the accuracies across the pairs that it is a part of. We report the accuracies of the merged models for the Same Label Space setting, and a linear probe accuracy on the representations of the merged model for the Different Label Space setting. * Note that here the merged models (bottom five) have the same size as the original, but the MoE and ensemble have a size twice the original.
@@ 5.2 Models fine-tuned with a shared label space
Method FLOPs Size Same Label Space Different Label Spaces
Clip Info Paint Real CUB Pets Dogs NABird
MoE* $\times1$ 2×2$\times$2 × 69.1 36.1 65.7 78.0 81.1 92.7 83.1 75.8
Ensemble* 2×2$\times$2 × 2×2$\times$2 × 63.6 30.3 61.0 74.7 80.5 92.8 82.1 76.1
Simple Avg 1×1$\times$1 × 1×1$\times$1 × 1.2 0.8 1.9 2.1 7.1 19.2 9.2 4.7
RegMean[14] 1×1$\times$1 × 1×1$\times$1 × 16.6 5.8 10.1 15.8 42.5 45.1 20.2 37.1
ZipIt[28] 1×1$\times$1 × 1×1$\times$1 × 26.9 12.2 27.1 37.4 67.5 83.6 60.0 56.3
Git Re-Basin[1] 1×1$\times$1 × 1×1$\times$1 × 18.2 7.8 18.8 26.5 66.2 80.2 62.6 59.4
PLeaS (Ours) 1×1$\times$1 × 1×1$\times$1 × 41.7 16.9 40.8 55.1 77.3 84.3 69.8 61.8
@@
We fine-tune ImageNet pre-trained ResNet-50 models on four different domains of the DomainNet[23] dataset: Clipart, Infograph, Painting and Real. We merge models trained (from different initializations) on different domains in a pair-wise fashion, and compute the accuracy of the merged model on both the domains. For each domain, we average the performance across all domain pairs. We report this in Table 1 for merged models at 1×1$\times$1 × size of the original model for different methods.
We also compute the average performance of each algorithm across all datasets at different size/FLOPs budgets, and plot this out in Fig. 3(a).
We find that PLeaS consistently outperforms ZipIt! at various FLOPs budgets. The gains are particularly striking for lower FLOP budgets, where PLeaS outperforms ZipIt! by up to 10%. The power of partial merging is also observed from these results, as one can see that increasing the flops by just 20% leads to massive improvements in the accuracies. Finally, one can see the importance of our layer-wise least squares, since it improves the performance of PLeaS over Permutation by over 20% on average at 1×1$\times$1 × model size. The gains decrease as we add more capacity to the merged model, which is expected.
5.3 Models with different label spaces
@@
We fine-tune models on CUB[31], NABirds[30], Oxford-IIIT Pets[21] and Stanford Dogs[16] datasets, and merge them upto the penultimate layer. Since the label spaces are different, we aim to evaluate the representations of the penultimate layer of these merged models by training a linear probe on top of the representations. We average the results in the same manner as for DomainNet, and report the performance of different methods for 1×1$\times$1 × model sizes in Tab 1. We also depict the compute-performance trade-off in Fig 3(b).
In App B.6, we follow the setting of [28], and use task specific heads from the original models to compute the accuracy of the merged model, which requires knowing the domain of the test data point.
Similar to the results on DomainNet, we observe that PLeaS has non-trivial gains over ZipIt! and Git Re-Basin, outperforming ZipIt! by 7% at 1×1$\times$1 × model size. These results also provide evidence of the effectiveness of our partial permutation scheme — Permutations can outperform ZipIt! at intermediate model budgets by up to 3%. A reason for this could be that the features of the models being merged are sufficiently different, leading to performance degradation if all of them are forced to be merged (a la ZipIt!). Our scheme retains some of these features in the intermediate layers, which could explain the better performance.
(a) Shared Label Space
(b) Different Label Spaces
@@
Figure 3: Memory-Performance trade-off for merged models: We merge pairs of models fine-tuned on different datasets, and compute the average performance across all four datasets for two settings - datasets with a Shared label space (Fig. 3(a)), and datasets with Different label spaces (Fig. 3(b)). We plot this average accuracy against the size of the final merged model, and find that PLeaS dominates the state-of-the-art methods.
@@ 5.4 Does PLeaS need data from the training domains?
Figure 4: Investigating the data requirement of PLeaS : We run PLeaS-Act and PLeaS using data from the actual domains or ImageNet (indicated by the suffix free) for both the Shared label space (Fig. 4(a)) and Different label spaces (Fig. 4(b)) settings. We plot the average accuracy (or linear probe accuracy) of each variant across all datasets against the relative size of the output model. We find minimal performance drops when using data from ImageNet instead of data from the original training domains.
To investigate the data requirements of our method, we compare the performance of PLeaS and PLeaS-free when merging models trained on different datasets. We also compare the effect of using activation matching to find the permutations for PLeaS , and we term this variant as PLeaS-Act . The performance-model size tradeoff is reported in Figs. 4(a) and 4(b) for the two settings of shared and different label spaces.
We find that PLeaS-free retains a similar performance when using ImageNet instead of the actual domain data for merging models on DomainNet, achieving a drop of less than 1% in accuracy at 1×1$\times$1 × model size. There is almost no drop at higher sizes of the merged model. We also find that PLeaS slightly outperforms PLeaS-Act , with the difference being around 1% at 1×1$\times$1 × model size. Notably, even on the more difficult task of merging models with different label spaces, using ImageNet data for computing activations can perform competitively to using the actual data: linear probing on the representations from PLeaS-free performs within 2% of the PLeaS at 1.2×1.2$\times$1.2 × model size, and the gap is less than 4% at 1×1$\times$1 × model size. This result is particularly encouraging, since it extends the practical applicability of PLeaS-free . Note that while we use data from the actual domains for linear probing, i.e. to assess the quality of the representations, we do not use it for actually merging the models.
We also find that PLeaS-Act performs similarly to PLeaS , with the latter performing better in the shared label space setting. We hypothesise that this is because we use only few batches of data to compute the activations for matching the two network, which leads to deteriorated permutations. However, the difference in performance decreases with increasing model size.
@@ 5.5 Merging models with the same initialization
Method Size Clipart Infograph Painting Real
TaskVectors 1.0 58.2 28.9 55.7 70.2
RegMean 1.0 58.9 29.0 57.4 71.8
ZipIt 1.0 54.1 25.2 51.2 66.1
PLeaS 1.0 60.1 29.6 58.0 72.0
PLeaS 1.2 64.2 31.9 61.8 76.0
Ensemble 2.0 64.4 32.0 62.0 76.1
In Tab 2, we evaluate the performance of our method for merging models fine-tuned from the same starting model. We compare against simple average (Task Vectors), ZipIt, and RegMean and find that the performance is similar across methods, with PLeaS being slightly better than the baselines. In fact, task vectors is an effective baseline here. However, we note that 20% extra parameters in the merged model can lead to closing the gap between the ensemble and the merged model produced by PLeaS , demonstrating the need for flexible merging methods.
6 Limitations and future work
The scope of this study is limited to merging models with the same architecture, and applying PLeaS to merge different architectures could be an interesting future direction. Since PLeaS is a two-stage algorithm, its running time is greater than some existing works [28, 12, 14]. However, since the second step can be computed in parallel for all layers, the runtime over-head is small. We further discuss this aspect in Section A.1. Finally, our study of model merging is limited to architectures with convolutions and residual connections, in line with prior work (e.g. [28, 33]). Extending this framework to other architectures such as transformers is another exciting future direction.

### v2

@@ 4.2 Permuted least squares
Suppose, for example, that the target merged model has the same architecture and size as each of the base models. Once the permutations, $P_{i}$, have been computed, we propose optimizing the weight matrices of the merged model by solving the following least-squares problem:
$W_{i}^{M}\in\smash{\operatorname*{arg\min}\limits_{W}}\|(Z_{i}^{A}+P_{i}Z_{i}^{B})W-(Z_{i+1}^{A}+P_{i+1}Z_{i+1}^{B})\|^{2},$
independently for each layer $i\in[L]$. This is motivated by the impressive performance of the ensemble method (e.g., [34] and Sec. 5), which retains two separate models and only averages the (permuted) activations at the last layer (pre-softmax): $\tilde{z}_{L+1}=z^{A}_{L+1}+z^{B}_{L+1}$. We aim to have our merged model approximate such activations. We inductively assume that the first $i-1$ layers are properly merged. Hence, the ensemble of the permuted features (of the $i$th layer) of the component models can be well approximated by the activations at the input of the $i$th layer of the merged model. We denote the ensembled features by $\tilde{Z}_{i}=(Z_{i}^{A}+P_{i}Z_{i}^{B})\in{\mathbb{R}}^{d_{i}\times n}$. The goal of the above optimization is to match the ensembled activation of the next layer, $\tilde{Z}_{i+1}=(Z_{i+1}^{A}+P_{i+1}Z_{i+1}^{B})$, with a linear transform of the input ensemble: $\tilde{Z}_{i}W$. We empirically validate this choice to use a permuted ensemble of features to optimize the weights of the merged model in Tab. 4 in Sec. B.1, where we compare with alternatives to Eq. 1.
This second step of PLeaS is similar to feature distillation. However, the key novelty arises from averaging the permuted features for transferring knowledge from multiple models. This is critical for accurate prediction. To show this, we compare PLeaS against RegMean [14], which optimizes an objective similar to Eq. 1 without the permutations and averaging, i.e. this method merges models by minimizing $\|Z_{i}^{A}W-Z_{i+1}^{A}\|^{2}+\|Z_{i}^{B}W-Z_{i+1}^{B}\|^{2}$. As we show in Sec. 5, RegMean performs poorly compared to PLeaS. Apart from the inference computation budget for the final model, PLeaS is completely hyperparameter free.
Note that the second step of PLeaS is fully compatible with the partial merging of Sec. 4.1 as well: we can directly set the values of $W_{i}^{M}$ corresponding to the unmerged features to be the respective values of $W_{i}^{A}$ and $W_{i}^{B}$.
While the objective in Eq. 1 can be minimized in closed form using Ordinary Least Squares (OLS), we practically implement it using gradient descent for ease of use with convolution layers. Given that the objective is convex if computed layer-wise, the weight matrices $W_{i}^{M}$ converge in relatively few (less than 100) steps of gradient descent. Further, we solve this optimization independently for each layer, so it can be efficiently parallelized.
@@ 4.3 Data requirements of PLeaS
PLeaS has two steps – the first step finds permutations to match features using weight or activation matching and the second step computes weight matrices to mimic the ensemble of the merged features more closely. In order to compute these features, one could use the data from the training domains, however, this may not be feasible for privacy or commercial reasons. Hence, we propose an alternative scheme–dubbed PLeaS-free–which uses a general vision dataset, like ImageNet, to compute the activations of the component models. These activations are then used to merge domain specific models without requiring any data from the training domains. In Sec. 5.4, we show that PLeaS-free suffers a minimal performance penalty compared to PLeaS, suggesting wider applicability in low/no data settings.
@@ 5 Experiments
We show the effectiveness of our method in merging models fine-tuned on different datasets with different initializations. We investigate the following research questions:
  1. 1.
    How does PLeaS  compare with prior work in merging models to produce a model of the same size (Sec. 5.2)?
  2. 2.
    What is the trade-off between size of the merged model and its performance for PLeaS  (Sec. 5.3)?
  3. 3.
    How does PLeaS  perform if one does not have access to the training data of the models being merged (Sec. 5.4)?
  4. 4.
    How does PLeaS  perform while merging models fine-tuned from the same initial model (Sec. 5.5), or different models trained on the same data (Sec. B.2)?
  5. 5.
    What is the impact of varying the objective in Eq. 1 on the performance of PLeaS (Sec. B.1)?
5.1 Experimental Setup
@@
Method FLOPs Memory Same Label Space Different Label Spaces
Clip Info Paint Real Avg CUB Pets Dogs NABird Avg
MoE* $\times1$ 2.1×2.1$\times$2.1 × 69.1 36.1 65.7 78.0 62.2 81.1 92.7 83.1 75.8 83.2
Ensemble* 2×2$\times$2 × 2×2$\times$2 × 63.6 30.3 61.0 74.7 57.4 80.5 92.8 82.1 76.1 82.9
Simple Avg [12] 1×1$\times$1 × 1×1$\times$1 × 1.2 0.8 1.9 2.1 1.5 7.1 19.2 9.2 4.7 10.1
RegMean [14] 1×1$\times$1 × 1×1$\times$1 × 16.6 5.8 10.1 15.8 12.1 42.5 45.1 20.2 37.1 36.2
ZipIt! [28] 1×1$\times$1 × 1×1$\times$1 × 26.9 12.2 27.1 37.4 25.9 67.5 83.6 60.0 56.3 66.9
Git Re-Basin [1] 1×1$\times$1 × 1×1$\times$1 × 18.2 7.8 18.8 26.5 17.8 66.2 80.2 62.6 59.4 67.1
MuDSC [32] 1×1$\times$1 × 1×1$\times$1 × 34.0 14.3 29.5 45.1 30.7 70.1 82.5 63.2 58.2 68.5
PLeaS (Ours) 1×1$\times$1 × 1×1$\times$1 × 41.7 16.9 40.8 55.1 38.6 75.2 85.0 69.6 69.7 74.9
Table 1: Merging pairs of models trained on different datasets: For each pair of datasets, we merge models and compute the final accuracy on the pair. To compute the final accuracy for a dataset, we average the accuracies across the pairs that the dataset is a part of. We report the accuracies of the merged models for the Same Label Space setting, and a linear probe accuracy on the representations of the merged model for the Different Label Space setting. * Note that here the merged models (bottom six) have the same size as the original, but the MoE and ensemble have a size twice the original.
@@
To obtain models for merging, we fine-tune ImageNet pre-trained ResNet models on other smaller datasets. We merge models trained (from different initializations) on different data domains in a pair-wise fashion, and compute the accuracy of the merged model on both the data domains. For each domain, we average the accuracy across all such pairs.
5.1.1 Datasets
Since we are dealing with classification models, we consider two sets of datasets (with shared and different label spaces) for training and merging models.
Datasets with a shared label space. We fine-tune ImageNet pre-trained ResNet-50 models on four different domains of the DomainNet [23] dataset: Clipart, Infograph, Painting and Real. These datasets share a label space with 345 classes and comprise of images in various styles.
@@ 5.1.1 Datasets
Datasets with different label spaces. We fine-tune models on CUB [31], NABirds [30], Oxford-IIIT Pets [21] and Stanford Dogs [16] datasets, and merge them up to the penultimate layer. Since the label spaces of these datasets are different, we aim to evaluate the representations of the penultimate layer of these merged models by training a linear probe on top of the representations. We average the results in the same manner as for DomainNet, and report the performance of different methods in Tab. 1. In Sec. B.4, we follow the setting of [28], using task specific heads from the original models to compute the accuracy of the merged model, which requires knowing the domain of each test data point.
5.1.2 Baselines
We compare our method against prior works including Git Re-Basin [1], Simple Averaging [12], RegMean [14], ZipIt! [28] and MuDSC [32]. Note that RegMean has similar data and compute requirements as PLeaS, and ZipIt! also supports partial merging of models like PLeaS. We also consider two practical upper bounds — training a router model based Mixture of Experts model (MoE), and ensembling the predictions (or activations) of the original models. The former requires storing both models and running one of them at inference in addition to the overhead of the router, hence having $\times1$ FLOPs and 2.1×2.1$\times$2.1 × memory requirements, while the latter requires running both the models in parallel, and hence has 2×2$\times$2 × FLOPs and memory requirements. We find that the performance of the ensemble and MoE models is close to the best performance of a single model on its training dataset.
For each task, we also report results for Permutations, which is the model obtained by weight averaging the component models after applying the permutations obtained from the first step of PLeaS. Following the recommendation of REPAIR [15], we recompute the batch-norm parameters of the model after merging for all methods. We run each merging experiment for three different seeds, and across two different initial models. We find that inter-run variation in performance is low, with the standard deviation usually being less than 1%. We report disaggregated results along with these standard deviations in Sec. B.3.
5.2 Merging for the same size
(a) ResNet-18 (Shared)
(b) ResNet-50 (Shared)
(c) ResNet-101 (Shared)
(d) ResNet-18 (Different)
(e) ResNet-50 (Different)
(f) ResNet-101 (Different)
@@ 5.2 Merging for the same size
Figure 3: Memory-Performance trade-off for merged models: We merge pairs of models fine-tuned on different datasets, and compute the average performance across all four datasets for two settings: datasets with a shared label space (top) and datasets with different label spaces (bottom). Plotting average accuracy against the final merged model size, we find that PLeaS dominates the state-of-the-art methods.
In Tab. 1, we report the domain-wise performance for merged models at 1×1$\times$1 × size of the original model for different methods. On DomainNet, we observe that PLeaS  outperforms the previous state-of-the-art method, MuDSC by over 8% on average. Similarly, On datasets with different labels spaces, PLeaS  is better than MuDSC by almost 6.5% on average. Further, PLeaS  vastly outperforms RegMean, a method which which performs Least Squares without permutations and has a similar computation overhead as PLeaS. This indicates that permuting the features before performing Least Squares is a crucial step which helps PLeaS  produce better models. Finally, These results also show the utility of the second step of PLeaS, as it boosts the performance over Git Re-Basin by a large amount, nearly doubling the accuracy on DomainNet.
5.3 Exploring the model size-accuracy tradeoff
We seek to demonstrate the effect of partial merging on the performance of the merged models. To do this, we merge models pairwise as above, and report the average of the performance of these merged models on their respective component domains for different sized ResNet models in Fig. 3.
We find that PLeaS consistently outperforms ZipIt! at various compute/memory budgets and for all model scales. The gains are particularly striking for lower memory budgets, where PLeaS outperforms ZipIt! by up to 10% for ResNet-50 ( Fig. 3(b)). The power of partial merging is also observed from these results, as one can see that increasing the flops by just 20% leads to massive improvements in the accuracies in the shared label settings. These results also provide evidence of the effectiveness of our partial permutation scheme — permutations can outperform ZipIt! at intermediate model budgets by up to 6% (e.g. for ResNet-101 with shared label spaces in Fig. 3(c)). We posit that this is because we can assign a non-uniform width multiplier across the layers of the merged model, which is important for larger models and those which are trained on disparate domains. As expected, the performance gap closes as the relative size of the merged models increases.
@@ 5.4 Does PLeaS need data from the training domains?
Figure 4: Investigating the data requirement of PLeaS : We run PLeaS and PLeaS-Weight using data from the actual domains or ImageNet (indicated by the suffix free) for both the Shared label space (Fig. 4(a)) and Different label spaces (Fig. 4(b)) settings for ResNet-50. We plot the average accuracy across all datasets against the relative size of the output model. We find minimal performance drops for PLeaS-free .
To investigate the data requirements of our method, we compare the performance of PLeaS and PLeaS-free when merging ResNet-50 models. We also compare the effect of using weight matching to find the permutations for PLeaS, and we term this variant as PLeaS-Weight. Note that this variant uses data only for the Least Squares step. The performance-model size tradeoff is reported in Figs. 4(a) and 4(b) for shared and different label spaces.
We find that PLeaS-free retains a similar performance when using ImageNet instead of the actual domain data for merging models on DomainNet, achieving a drop of less than 1% in accuracy at 1×1$\times$1 × model size. There is almost no drop at higher sizes of the merged model. Notably, even on the more difficult task of merging models with different label spaces, using ImageNet data for computing activations can perform competitively to using the actual data: linear probing on the representations from PLeaS-free performs within 2% of the PLeaS at 1.2×1.2$\times$1.2 × model size, and the gap is less than 4% at 1×1$\times$1 × model size. This result is particularly encouraging, since it extends the practical applicability of PLeaS-free to scenarios where data from the training domains may not be available. Note that while we must use data from the actual domains for linear probing, i.e. to assess the quality of the representations, we do not use it for actually merging the models. We also find that PLeaS-Weight performs similarly to PLeaS  in the both the shared and different label space settings for ResNet-50. Further, PLeaS-Weight is less affected when ImageNet data is used, since the permutations computed by PLeaS-Weight do not depend on the data.
@@ 5.5 Merging models with the same initialization
Method Size Clipart Infograph Painting Real
Simple Avg 1.0 58.2 28.9 55.7 70.2
RegMean 1.0 58.9 29.0 57.4 71.8
MuDSC 1.0 57.8 55.8 55.1 67.6
PLeaS 1.0 59.6 29.5 58.0 72.0
PLeaS 1.2 64.2 31.9 61.8 75.9
Ensemble 2.0 64.4 32.0 62.0 76.1
@@
In Tab 2, we evaluate the performance of our method for merging ResNet-50 models fine-tuned from the same starting model. We compare against simple average (Task Vectors), MuDSC, and RegMean and find that the performance is similar across methods, with PLeaS being slightly better than the baselines. In fact, task vectors is an effective baseline here. However, we note that 20% extra parameters in the merged model can lead to closing the gap between the ensemble and the merged model produced by PLeaS, demonstrating the need for flexible merging methods.
5.5.1 Merging ViTs
Method CIFAR-50+50 Other Datasets
Joint Avg NABirds CUB Pets Dogs
Simple Avg 72.6 84.5 7.8 65.9 86.1 60.4
RegMean 72.7 84.7 7.7 66.2 85.1 58.5
MuDSC 72.8 84.9 8.0 66.1 86.1 60.6
PLeaS 73.3 85.1 8.3 66.7 86.5 61.6
Table 3: Merging ViT models: We merge pairs of ViT models fine-tuned from the same initialization and report the performance. PLeaS  can outperform baselines across datasets.
In Tab. 3, we present results of merging CLIP style ViT models. In accordance with prior work [32], we merge models starting from the same initialization. We consider two settings – the different label space setting described in Sec. 5.1 and CIFAR50+50 from [28, 32]. In the latter, the 100 labels from CIFAR-100 are partitioned into 2 sets of 50 labels each, and a ViT is trained on each of these sets (using a CLIP-like loss). More details on the setup are described in Appendix A. The accuracy of the merged model is reported on both these partitions separately by considering only 50 classes at a time (denoted by Avg in the table), as well as on the Joint CIFAR-100 dataset (by considering all 100 classes together). Note that since we use CLIP-like models, we can use the language head directly for classification despite different label spaces. For the other datasets, we follow the protocol from Sec. 5.1.
We observe that PLeaS  boosts the performance slightly over the baselines. For example, it increases performance by 0.7% on CIFAR-100 over Task Vectors. PLeaS  also out-performs MuDSC by around 0.6% on the larger datasets. While these gains are small, they are non-trivial and are more than the boost reported by the previous state-of-the-art method (MuDSC). We believe this presents an exciting opportunity to further study the symmetries and invariances in transformers in order to merge them better.
6 Limitations
The scope of this study is limited to merging models with the same architecture, and applying PLeaS to merge different architectures could be an interesting future direction. Since PLeaS is a two-stage algorithm, its running time is greater than some existing works [28, 12, 14]. However, since the second step can be computed in parallel for all layers, as discussed in Sec. A.1.